OpenStudy (anonymous):

Solve the equation .Check for extraneous solutions 2|4-3x|=4x+6

OpenStudy (shubhamsrg):

you have this' |4 - 3x| = 2x + 3 you'll get 2 eqns from here.. 4-3x = 2x+3 and 4-3x = -(2x+3) solve for x from both its just like this |x| =1 then x= 1 and -1..

OpenStudy (anonymous):

Thank you.